ian spent 3/5 of his money on a book and 1/4 of the remainder on a pen.If a book cost 350 more than the pen how much money did he have at first.?
Answer by TimothyLamb(4379)

b = (3/5)M
---
r = M - b
r = M - (3/5)M
r = (5/5)M - (3/5)M
r = (2/5)M
---
p = (1/4)r
p = (1/4)(2/5)M
p = (1/10)M
---
b = p + 350
---
b = p + 350
(3/5)M = (1/4)r + 350
(3/5)M = (1/10)M + 350
(6/10)M - (1/10)M = 350
(5/10)M = 350
(1/2)M = 350
M = 700
---
answer: he had 700
NOTE:
money units are not given in the problem statement.
money units can be anything such as dollar, euro, peso, etc.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php